# Auditing and Generating Synthetic Data with Controllable Trust Trade-offs

### v3

@@ Methods
  • Robustness. Robustness refers to the accuracy of a predictive model trained on synthetic data and evaluated on real test points in the presence of imperceptible, worst-case adversarial perturbations. We use black box, greedy attacks on utility classifiers for tabular and time-series, as in [70, 71, 72, 72, 73, 74, 75] (See Supplementary Information C.5).
@@
    Evaluate: Given the specific synthetic data application and relevant policies and regulations, the governance expert and data science board collaboratively establish multiple quantitative metrics for each trust dimension. These metrics serve to evaluate the synthetic data’s adherence to the requirements necessary for meeting socio-technical safeguards within each dimension. In Extended Data Table 4, a comprehensive set of metrics is presented for each dimension, chosen to strike a balance between interpretability and risk assessment. It’s important to note that while these metrics represent each dimension and modality, they are not exhaustive. Our framework can seamlessly accommodate additional specialized metrics as needed. These metrics are then assessed by the data scientist.
@@
    Aggregate: The interpretation and communication of these metrics within each dimension pose a significant challenge. Dealing with numerous metrics with different polarities and dynamic ranges can be overwhelming for internal reviewers and regulators. In social sciences, it’s common to aggregate metrics into a single score or index [76]. Indices serve as powerful tools for simplifying complex information into an accessible format that can be interpreted and understood by a wide range of stakeholders. They facilitate straightforward communication, enabling easy comparisons and benchmarking. In our Methods Section, we address the issues related to varying ranges and polarities and aggregate the metrics within each dimension into a trust dimension index that falls within the range of 0 to 1, where 0 signifies poor conformity with the dimension requirements, and 1 indicates high conformity. This trust dimension index can be interpreted as a measure of compliance probability. Our aggregation method relies on the copula technique [77] which provides us with this intuitive probabilistic interpretation.
@@
    Re-weight: Each synthetic dataset is now represented by trust dimension indices, denoted as $\pi_{T}$, where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness. Considering the specific application, associated policies, and desired trade-offs between trust dimensions, internal reviewers collaborate with governance experts and the application owner to establish the trustworthiness profile. This profile is expressed through tradeoff weights, symbolized as $\omega_{T}$, which indicate the relative importance of the trust indices necessary to fulfill performance and policy requirements. For instance, when training downstream tasks on-site, privacy may not be a necessity for the synthetic data, but it becomes essential for the predictive model. However, when conducting training in a public cloud environment, ensuring the privacy of synthetic data is mandatory. These varying requirements lead to different trustworthiness profiles with distinct trade-offs between privacy and utility for the synthetic data. Detailed examples of other possible trustworthiness profiles can be found in Extended Data Table 2.
@@
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \tau_{\text{Trust}}(\omega)=\exp\left(\sum_{T}\omega_{T}\log(\pi_{T})\right),$ (3)
@@
    $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R^{\alpha}_{\text{trust}}=\log(\overline{\tau_{\text{Trust}}}(\omega))-\alpha\log(\Delta_{\tau}(\omega)).$ (4)
@@
    Report: To promote transparency and accountability, our framework defines templates for communicating auditing results in the form of an audit report. An example of the audit report is given in Supplementary Information G. The audit report can be submitted to a regulator or an external third-party certifier that probe the validity of the conclusions of the internal audit report.
@@
Use cases and Experimental Setups
@@
In our auditing setup, real data serves as the baseline, and we emphasize the contrast with synthetic generated data. The real data is split into training ($D_{r,\text{train}}$), validation ($D_{r,\text{val}}$), and test ($D_{r,\text{test}}$) sets through random sampling without replacement, repeated five times with different seeds, creating five real data folds ($\mathcal{D}_{r}$). Each generative AI method listed in Extended Data Table 3 is trained five times independently on each real data fold. For TrustFormers, both non-private (TF(n-p)) and private (TF(p-$\varepsilon=1$) or TF(p-$\varepsilon=3$)) versions are trained. Trustworthiness Index Driven Model selection is performed independently within each fold, leading to the selection of a checkpoint $t^{*}$. Synthetic datasets are generated by sampling from each generator trained on a specific fold. This results in $\mathcal{D}_{s}=\{D^{\ell}_{s}\}_{\ell=1}^{5}$. Audits on downstream tasks, assess utility, fairness, and robustness evaluated on the corresponding real test data folds. A non-private tabular RoBERTa Embedding (E) [3], trained on the real data $D_{r}$ (excluding task labels), is used for certain metrics in Extended Data Table 4.
@@
Time Series Use Cases
Time-Series Use Case I: MIMIC-III Controllable Trust trade-offs on Healthcare Data
@@
Extended Data Table 5 summarizes the trust dimension indices of selected TrustFormers models, where the downstream tasks are evaluated on the real test set. Interestingly similar to tabular data, we see that TrustFormer synthetic data outperforms real data on all trust dimensions. Interestingly the fairness index of real data is the lowest, and synthetic data therefore improves the fairness/utility tradeoff . Similar observations on the relationship between trust constraints and trust trade-offs we made on tabular data hold for the time series case.
Time Series Use Case II: Fraud Detection, Deep Dive on Utility and Privacy trade-offs in Synthetic Data
Training RoBERTa-like Embedding
To train TabRoBERTa on our transaction dataset, we constructed samples as sliding windows of 10 transactions, using a stride of 5. We excluded the label column, "isFraud?", during training to prevent biasing the learned representation for the downstream fraud detection task. We masked $15\%$ of a sample’s fields, replacing them with the [MASK] token, and predicted the original field token using cross-entropy loss. We used DP-SGD for transformer models [31] to train various RoBERTa-like models with differing degrees of privacy, ranging from highly private ($\varepsilon=1$) to non-private ($\varepsilon=1000$). Additionally, we trained a RoBERTa model without private training (see the columns labeled "Private" and "Non-Private" in Table 1).
Synthetic data generation
We generated several privacy-preserving synthetic datasets using our non-private pretrained TabGPT model. For model selection in this experiment we relied on a fidelity validation of the TabGPT model. To generate private synthetic data, we used a private sampling technique [85] , which involves adding Laplacian noise with controlled variance (dependent on the user-provided $\varepsilon$ value) to the probability distribution over the generated tokens from the non-private GPT model. This is a form of output-perturbation methods that guarantees differential privacy.We generated seven datasets with varying privacy levels, from highly private ($\varepsilon=0.1$) to non-private ($\varepsilon=200$), as shown in the rows labeled "Synthetic" in Table 1. Additionally, we considered real card transaction data and synthetically generated data without private sampling. Refer to Section A.5 for more details on private sampling.
Training the downstream Fraud Detection Model
Ablation Studies
Debiasing downstream tasks trained on Synthetic Data
We notice on tabular datasets that while private synthetic data with high performing robustness index are still competitive in terms of their utility, the ones with high fairness index are not competitive in terms of their utility. In order to further improve the fairness, privacy and utility tradeoffs, we explore including debiased downstream classifiers in Supplementary Information Tables 6, 10 and 14. Debiasing leads to an improvement fairness while maintaining utility of private synthetic data.
Independent Copula Versus Dependent Copula
In our aggregation we used an independent copula as a mean to aggregate the metrics. We explore here using a Gaussian copula to capture the dependencies between metrics and see how it impacts the overall ranking of the synthetic datasets on all tabular datasets we consider. We explore Gaussian copulas, since estimating dependent copulas from data is cursed in high dimension. To estimate the Gaussian copula, we compute the mean and covariance matrix for the concatenation of all metrics under a trust dimension, and compute the multivariate Gaussian CDF. This comes with two caveats: 1) the dependency may not be Gaussian 2) given small number of samples the covariance matrix is not well estimated and is low rank and inverting it in the computation of the multivariate CDF leads to numerical instabilities that we circumvent with a regularization of the covariance matrix $\Sigma$ by $\Sigma+\lambda I$, for $\lambda=10^{-3}$. We first report the average time in seconds to compute the independent copula on tabular use cases on an Intel (R) Xeon (R) CPU E5-2667 v2 @ 3.30GHz that is of $11s$ versus $1700s$ for the Gaussian copula (See Supplementary Information E Table 3). Second we report the overlap between the ranked list of synthetic data returned by the independent copula $L_{IC}$ and the one returned by the Gaussian copula $L_{GC}$:
[…]
From Supplementary Information E Figure 4 we see that there is 60-70% overlap between the top-10 ranking synthetic datasets returned by independent and Gaussian copula. In light of these results, given the simplicity and the speed of the independent copula and the numerous issues of the Gaussian copula in terms of its slow computation and bad conditioning and unfavorable statistical guarantees, the independent copula is preferable and offers a good computational tradeoff in estimating the copula.
Membership attacks
The membership attacks we used in our privacy auditing were based on nearest neighbor attacks that assess memorization issues in synthetic data. Identity disclosure attacks assess privacy in synthetic data beyond memorization [86],[87]. We explore here the impact of other targeted membership attacks based on likelihood ratios between synthetic data and a reference to estimate the membership [88]. This attack is the state of the art membership inference method. We train this attack on the Bank Marketing dataset, evaluating synthetic data of three baselines: the Gaussian copula, CTGAN and PateGAN. We give in in Extended Table 7, $1-\textrm{AUC}$ of this attack, where the Area Under the Curve (AUC) is evaluated on the real data with members and non members of the training set. Higher values of $1-\textrm{AUC}$ mean that the membership attack is less successful at identifying members of the real data, meaning higher privacy. We see in Extended Table 7 that this membership attack lead to the same ranking of synthetic datasets as our privacy index (computed without the inclusion of the membership attack). Those membership inference attacks can be seamlessly added to our privacy index computation.
Iterative Collapse Risk of Synthetic Data
As synthetic data becomes more and more prevalent, the risks of iteratively training generative models on synthetic data becomes more and more pressing. Authors in [89] referred to this issue as the iterative curse or iterative collapse in synthetic data. They also showed that the fidelity of the synthetic data deteriorates as we iteratively train generative models on synthetic data. We show in Extended Figure 3 on the Bank Makerting dataset, the audits of synthetic data obtained from training models successively on synthetic data generated from the previous iteration. As we iteratively use synthetic data to train generative models, we see in Extended Data Figure 3 that while the fidelity index of the synthetic datasets collapses, the privacy index improves as expected. The utility of the synthetic data on downstream tasks fluctuates but is essentially constant. Note that the utility of these models does not deteriorate, since the utility of these synthetic data at first iteration is not high.
Data availability
Data used in the paper is available online and open source. See Extended Data Table 1 for references.
Code availability
1 Figures and Tables
Figure 1:
Figure 2:
Figure 3:
@@ 1 Figures and Tables
Table 1:
@@
Table 2:
@@
Table 3:
2 Figure and Table Captions
Caption of Figure 1 Summary diagram of our proposed holistic synthetic data auditing framework. For each trust dimension (fidelity, privacy, utility, fairness, and robustness), we evaluate multiple metrics on the synthetic data and quantify their uncertainty. Metrics are aggregated within each trust dimension, which results in trust dimension indices. These indices are re-weighted with desired trust trade-offs to produce the trustworthiness index. Different synthetic datasets are then ranked using this trustworthiness index, and a summary of the audit is written to an audit report. The ranking produced by our audit enables comparison of different synthetic data produced by various generative modeling techniques, and aids the model selection process for a given generation technique, allowing its alignment with prescribed safeguards. The model selection is performed via trustworthiness index driven cross-validation, which results in controllable trust trade-offs by producing new ranks for different desired weighing trade-offs for a given application and use case.
Caption of Table 1 Performance (F1-macro) of the fraud classifier on the test set of credit card transactions for different training choices of classifier (rows) and TabRoBERTa features extractor (columns).
Caption of Table 2 Trust Indices of $\text{BioGPT}_{\text{Finetuned}}$ on MIMIC-III notes. Robustness dimension is not evaluated herein. trustworthiness index Ranking of synthetic data sampled from different epoch during the finetuning of BioGPT model (Note that the robustness dimension was not considered) . The ranking corresponds to different trust tradeoffs $\omega$: for $w_{U}$ that is accuracy driven, we see that the last epoch is outperforming the other ones; When in addition we consider the fairness of the prediction ($\omega_{UF}$) the last epoch (epoch 9) ranks third and the epoch 7 presents better utility/ Fairness trade-offs.
Caption of Figure 2 Summary of auditing and ranking results on the Bank Marketing dataset using the trustworthiness index given in (4) for $\alpha=0$. (a) and (b) show trust dimension indices $\pi_{T}$( where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness), and their "variance" ($\Delta_{T}$) on TrustFormer (TF) and baseline models. The format is $\pi_{T}(\Delta_{T})||$ Name of the synthetic data model. (c) shows the ranking of the models across different trustworthiness profiles $\omega$ given in Extended Table 2.
Caption of Figure 3 Summary of auditing and ranking results on the Bank Marketing dataset using the trustworthiness index given in (4) for $\alpha=0.1$. This means that the uncertainty is taking into account in both model selection and in ranking the synthetic data using the trustworthiness index. (a) and (b) show trust dimension indices $\pi_{T}$( where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness), and their "variance" ($\Delta_{T}$) on TrustFormer (TF) and baseline models. The format is $\pi_{T}(\Delta_{T})||$ Name of the synthetic data model. (c) shows the ranking of the models across different trustworthiness profiles $\omega$ given in Extended Table 2.

### v4

@@ Methods
  • Robustness. Robustness refers to the accuracy of a predictive model trained on synthetic data and evaluated on real test points in the presence of imperceptible, worst-case adversarial perturbations. We use black box, greedy attacks on utility classifiers for tabular and time-series, as in [14, 25, 112, 112, 66, 73, 4] (see Supplementary Information -P).
@@
    Evaluate: Given the specific synthetic data application and relevant policies and regulations, the governance expert and data science board collaboratively establish multiple quantitative metrics for each trust dimension. These metrics serve to evaluate the synthetic data’s adherence to the requirements necessary for meeting socio-technical safeguards within each dimension. In Table I, a comprehensive set of metrics is presented for each dimension, chosen to strike a balance between interpretability and risk assessment. It’s important to note that while these metrics represent each dimension and modality, they are not exhaustive. Our framework can seamlessly accommodate additional specialized metrics as needed. These metrics are then assessed by the data scientist.
@@
    Aggregate: The interpretation and communication of these metrics within each dimension pose a significant challenge. Dealing with numerous metrics with different polarities and dynamic ranges can be overwhelming for internal reviewers and regulators. In social sciences, it’s common to aggregate metrics into a single score or index [45]. Indices serve as powerful tools for simplifying complex information into an accessible format that can be interpreted and understood by a wide range of stakeholders. They facilitate straightforward communication, enabling easy comparisons and benchmarking. We address the issues related to varying ranges and polarities and aggregate the metrics within each dimension into a trust dimension index that falls within the range of 0 to 1, where 0 signifies poor conformity with the dimension requirements, and 1 indicates high conformity. This trust dimension index can be interpreted as a measure of compliance probability. Our aggregation method relies on the copula technique [105] which provides us with this intuitive probabilistic interpretation. In Supplementary Information -A we explain the copula method that consists in normalizing metrics under trust dimension using global CDFs estimated across synthetic data, and followed by a geometric mean.
@@
    Re-weight: Each synthetic dataset is now represented by trust dimension indices, denoted as $\pi_{T}$, where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness. Considering the specific application, associated policies, and desired trade-offs between trust dimensions, internal reviewers collaborate with governance experts and the application owner to establish the trustworthiness profile. This profile is expressed through tradeoff weights, symbolized as $\omega_{T}$, which indicate the relative importance of the trust indices necessary to fulfill performance and policy requirements. For instance, when training downstream tasks on-site, privacy may not be a necessity for the synthetic data, but it becomes essential for the predictive model. However, when conducting training in a public cloud environment, ensuring the privacy of synthetic data is mandatory. These varying requirements lead to different trustworthiness profiles with distinct trade-offs between privacy and utility for the synthetic data. Detailed examples of other possible trustworthiness profiles can be found in Table II.
@@
    $\tau_{\text{Trust}}(\omega)=\exp\left(\sum_{T}\omega_{T}\log(\pi_{T})\right),$ (3)
@@
    $R^{\alpha}_{\text{trust}}=\log(\overline{\tau_{\text{Trust}}}(\omega))-\alpha\log(\Delta_{\tau}(\omega)).$ (2)
@@
    Report: To promote transparency and accountability, our framework defines templates for communicating auditing results in the form of an audit report. An example of the audit report is given in Supplementary Information -T. The audit report can be submitted to a regulator or an external third-party certifier that probe the validity of the conclusions of the internal audit report.
@@
Generative AI Method Trust Constraint
Non-Private & Private TrustFormers
(Conditional) TrustFormer GPT (non-private): TF($\omega$,n-p) Fidelity/Utility
(Conditional) TrustFormer GPT Trained with Differential Private SGD : TF($\omega$,p-$\varepsilon$) Fidelity/Privacy Preserving synthetic Data/Utility
Differential Private Sampling From non-private (Conditional)TrustFormer [72, 92] Fidelity/ Privacy Preserving synthetic Data/utility
Non-private Baselines
Gaussian Copula [84] Gaussian Copula (n-p) Fidelity
Conditional Tabular GAN[84]CTGAN(np) Fidelity/utility
Private Baselines
Differential Private Probabilistic Graphical Model [74] DP-PGM(p-$\varepsilon$) Fidelity/Privacy Preserving synthetic Data
DP-PGM (targeted)[74] DP-PGM(target,p-$\varepsilon$) Fidelity/Privacy Preserving synthetic Data/Utility
(Conditional) Differential Private-GAN [111, 88] DP-GAN(p-$\varepsilon$) Fidelity/ Privacy Preserving synthetic Data/Utility
(Conditional) PATE-GAN [113, 88]PATE-GAN(p-$\varepsilon$) Fidelity/Privacy Preserving synthetic Data/Utility
TABLE IV: Generative Models and TrustFormer Models audited within our framework. For private models we consider privacy budgets $\varepsilon=1$ or $3$.
In the following sections we present the training and auditing of various data generation techniques, including our TrustFormer Models, across a range of use cases encompassing tabular, time series, and natural language data (refer Table III). We also conduct audits on synthetic data generated by state-of-the-art diffusion models, particularly in the computer vision domain, focusing on the ImageNet classification task and assessing their generalization under distribution shifts. The models under audit are summarized in Table IV, and we employ our trust dimension indices to assess each model’s compliance with specific trust dimensions. For a given trustworthiness profile weights $\omega$ (refer to Table II for examples and their interpretation), we rely on our trustworthiness index to conduct TrustFormer model selection and rank the models based on their alignment with predefined safeguards (given in the trustworthiness profile $\omega$).
@@
In this Section, we showcase our auditing framework and trustworthiness index driven model selection for TrustFormer on three tabular datasets: Bank Marketing Dataset[79], Recruitment Dataset [41], and the Law School Admission Council Dataset [110]. Please refer to Table III for details of the datasets, downstream tasks , safeguards and TF models training details.
@@
In our auditing setup, real data serves as the baseline, and we emphasize the contrast with synthetic generated data. The real data is split into training ($D_{r,\text{train}}$), validation ($D_{r,\text{val}}$), and test ($D_{r,\text{test}}$) sets through random sampling without replacement, repeated five times with different seeds, creating five real data folds ($\mathcal{D}_{r}$). Each generative AI method listed in Table IV is trained five times independently on each real data fold. For TrustFormers, both non-private (TF(n-p)) and private (TF(p-$\varepsilon=1$) or TF(p-$\varepsilon=3$)) versions are trained. Trustworthiness Index Driven Model selection is performed independently within each fold, leading to the selection of a checkpoint $t^{*}$. Synthetic datasets are generated by sampling from each generator trained on a specific fold. This results in $\mathcal{D}_{s}=\{D^{\ell}_{s}\}_{\ell=1}^{5}$. Audits on downstream tasks, assess utility, fairness, and robustness evaluated on the corresponding real test data folds. A non-private tabular RoBERTa Embedding (E) [82], trained on the real data $D_{r}$ (excluding task labels), is used for certain metrics in Table I.
@@
Figure 3: Summary of auditing and ranking results on the Bank Marketing dataset using the trustworthiness index given in (2) for $\alpha=0$. (a) and (b) show trust dimension indices $\pi_{T}$( where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness), and their "variance" ($\Delta_{T}$) on TrustFormer (TF) and baseline models. The format is $\pi_{T}(\Delta_{T})||$ Name of the synthetic data model. (c) shows the ranking of the models across different trustworthiness profiles $\omega$ given in Table II.
Figure 4: Summary of auditing and ranking results on the Bank Marketing dataset using the trustworthiness index given in (2) for $\alpha=0.1$. This means that the uncertainty is taking into account in both model selection and in ranking the synthetic data using the trustworthiness index. (a) and (b) show trust dimension indices $\pi_{T}$ (where ‘T’ corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness), and their “variance” ($\Delta_{T}$) on TrustFormer (TF) and baseline models. The format is $\pi_{T}(\Delta_{T})||$ Name of the synthetic data model. (c) shows the ranking of the models across different trustworthiness profiles $\omega$ given in Table II.
Figure 5: Summary of auditing and ranking results on the recruitment dataset using the trustworthiness index given in (2) for $\alpha=0$. (a) and (b) show trust dimension indices $\pi_{T}$ (where "T" corresponds to Fidelity, Privacy, Utility, Fairness, or Robustness), and their "variance" ($\Delta_{T}$) on TrustFormer (TF) and baseline models. The format is $\pi_{T}(\Delta_{T})||$ Name of the synthetic data model. (c) shows the ranking of the models across different trustworthiness profiles $\omega$ given in Extended Table II.
Audit Results
For the Bank Marketing Dataset, we present average trust dimension indices and their "variances" for non-private and private synthetic data generated by TrustFormers models (TF) in Figure 3 (a) and (b). In this case, TF models were selected across various trustworthiness profiles (as outlined in Table II) using the trustworthiness index from Equation (2) for $\alpha=0$. We also include results for non-private and private baselines (refer to Table IV), shown in panels (a) and (b). Additionally, we provide information about the utility, fairness, and robustness of downstream tasks trained on real data as a reference for both panels (a) and (b). Panel (c) displays the ranking according to trustworthiness index of synthetic data across different trustworthiness profiles based on the same $\alpha$ value. Similar plots for the Recruitment and Law School datasets can be found in Figure 5 and in Supplementary Information -V Figure 10, respectively, for $\alpha=0$. To account for data split uncertainties, we’ve included audit results for $\alpha=0.1$, depicted in Figure 4 for the Law School dataset, and in the Supplementary Information for the other datasets (see Figures 11 and 12 in Supplementary Information -V).
For both Bank Marketing dataset and Recruitment datasets we see that in both cases $\alpha=0$ and $0.1$, TF synthetic data outperforms other baselines across all trustworthiness profiles $\omega$ in terms of its trustworthiness index . We see in Panel (c) Figure 3 and Figure 5, a clear phase transition between private and non private TF models depending on wether the trustworthiness profile highlights privacy as a requirement or not. This effect is achieved through our trustworthiness index model selection that offers control over trust trade-offs. On the other hand, for the law school dataset, TF models fall short with respect to baselines when uncertainty is not considered ($\alpha=0$, Supplementary Figure 10), but lead to top performing synthetic data when uncertainty is considered ($\alpha=0.1$, Figure 4). This highlights the importance of assessing the uncertainty in auditing synthetic data.
Model Fidelity Privacy Utility Fairness Robustness
Non-private TrustFormer
TF ($\omega_{UR}$, n-p, mn) 0.70 0.11 0.57 0.40 0.55
TF ($\omega_{e(PUF)}$,$\omega_{U}$,$\omega_{UF}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$,$\omega_{UFR}$, n-p, mn) 0.37 0.36 0.67 0.42 0.50
TF ($\omega_{e(PU)}$,$\omega_{PUR}$, n-p, mn) 0.53 0.21 0.51 0.29 0.49
TF ($\omega_{PU}$, n-p, mn) 0.81 0.42 0.56 0.25 0.39
TF ($\omega_{UR}$, n-p, topk) 0.70 0.21 0.52 0.43 0.60
TF ($\omega_{UF}$,$\omega_{UFR}$, n-p, topk) 0.86 0.49 0.61 0.44 0.52
TF ($\omega_{U}$, n-p, topk) 0.62 0.24 0.62 0.32 0.69
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{PU}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$,$\omega_{PUR}$, n-p, topk) 0.77 0.54 0.60 0.29 0.64
Private TrustFormer
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{PU}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$, p - ϵ=italic-$\epsilon=$3, mn) 0.34 0.93 0.30 0.58 0.41
TF ($\omega_{U}$,$\omega_{UF}$,$\omega_{UFR}$,$\omega_{UR}$,$\omega_{PUR}$, p - ϵ=italic-$\epsilon=$3, mn) 0.19 0.69 0.44 0.55 0.47
TF ($\omega_{PU}$, p - ϵ=italic-$\epsilon=$3, topk) 0.15 1.00 0.22 0.83 0.46
TF ($\omega_{PUR}$, p - ϵ=italic-$\epsilon=$3, topk) 0.35 0.89 0.40 0.46 0.50
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{UF}$,$\omega_{all}$, p - ϵ=italic-$\epsilon=$3, topk) 0.40 0.81 0.33 0.57 0.29
TF ($\omega_{e(UF)r(R)}$, p - ϵ=italic-$\epsilon=$3, topk) 0.34 0.75 0.26 0.31 0.40
TF ($\omega_{UFR}$,$\omega_{UR}$, p - ϵ=italic-$\epsilon=$3, topk) 0.42 0.62 0.41 0.60 0.40
TF ($\omega_{U}$, p - ϵ=italic-$\epsilon=$3, topk) 0.64 0.12 0.41 0.59 0.34
Real Data N/A N/A 0.44 0.07 0.67
Table V: MIMIC-III/ In-Hospital Mortality downstream task evaluation: trust dimension indices of TrustFormer models. In bold highest index within each group of synthetic data. In blue highest value across all methods including real data.
Non-Private TrustFormer
TF ($\omega_{UR}$, n-p, mn) 13 15 15 14 15 15 5 6 5 6
TF ($\omega_{e(PUF)}$,$\omega_{U}$,$\omega_{UF}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$,$\omega_{UFR}$, n-p, mn) 8 7 9 5 8 5 1 1 1 3
TF ($\omega_{e(PU)}$,$\omega_{PUR}$, n-p, mn) 15 14 14 15 14 14 8 14 12 7
TF ($\omega_{PU}$, n-p, mn) 9 9 10 10 9 10 6 15 14 8
TF ($\omega_{UR}$, n-p, topk) 7 12 12 9 13 13 7 7 4 4
TF ($\omega_{UF}$,$\omega_{UFR}$, n-p, topk) 1 1 1 1 4 3 3 2 2 5
TF ($\omega_{U}$, n-p, topk) 6 10 11 11 12 7 2 8 3 1
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{PU}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$,$\omega_{PUR}$, n-p, topk) 2 2 3 3 2 1 4 12 7 2
Private TrustFormer
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{PU}$,$\omega_{e(UF)r(R)}$,$\omega_{all}$, p - ϵ=italic-$\epsilon=$3, mn) 5 5 5 7 5 6 14 13 13 13
TF ($\omega_{U}$,$\omega_{UF}$,$\omega_{UFR}$,$\omega_{UR}$,$\omega_{PUR}$, p - ϵ=italic-$\epsilon=$3, mn) 11 8 7 8 3 4 9 4 6 9
TF ($\omega_{PU}$, p - ϵ=italic-$\epsilon=$3, topk) 12 11 8 12 10 9 16 10 10 15
TF ($\omega_{PUR}$, p - ϵ=italic-$\epsilon=$3, topk) 3 3 2 6 1 2 12 11 9 10
TF ($\omega_{e(PU)}$,$\omega_{e(PUF)}$,$\omega_{UF}$,$\omega_{all}$, p - ϵ=italic-$\epsilon=$3, topk) 10 6 6 4 6 11 13 9 15 16
TF ($\omega_{e(UF)r(R)}$, p - ϵ=italic-$\epsilon=$3, topk) 14 13 13 16 11 12 15 16 16 14
TF ($\omega_{UFR}$,$\omega_{UR}$, p - ϵ=italic-$\epsilon=$3, topk) 4 4 4 2 7 8 10 3 8 11
TF ($\omega_{U}$, p - ϵ=italic-$\epsilon=$3, topk) 16 16 16 13 16 16 11 5 11 12
Table VI: MIMIC-III synthetic dataset ranking using the trustworthiness index corresponding to the trade-off weight $\omega$. We see that two models stand out across different trade-offs and they correspond to different decoding strategies.
Time-Series Use Cases
@@
Table V summarizes the trust dimension indices of selected TrustFormers models, where the downstream tasks are evaluated on the real test set. Interestingly similar to tabular data, we see that TrustFormer synthetic data outperforms real data on all trust dimensions. Interestingly the fairness index of real data is the lowest, and synthetic data therefore improves the fairness/utility tradeoff . Similar observations on the relationship between trust constraints and trust trade-offs we made on tabular data hold for the time series case.
Analysis of Results
We compare the performance of the two decoding strategies considered (multinomial and top-k decoding) for synthetic data generation from TF models and study how it impact trust trade-offs. Note that we used herein fixed data splits from the literature and we don’t report therefore the uncertainty of the audit. Table VI gives the ranking of these synthetic datasets using the trustworthiness index for all trustworthiness profiles. We see that for some trustworthiness profiles, non-private TrustFormer with multinomial decoding stands out, while top-k decoding outperforms it for other profiles. On private models we see that TF models with top-k decoding are the best at balancing the privacy/utility trade-offs. This shows that our auditing framework highlights the effect of hyper-parameters choices such as decoding strategies and their impacts on all trust dimensions
@@
TABLE VII: Performance (F1-macro) of the fraud classifier on the test set of credit card transactions for different training choices of classifier (rows) and TabRoBERTa features extractor (columns).
@@
TABLE VIII: Trust Indices of $\text{BioGPT}_{\text{Finetuned}}$ on MIMIC-III notes. Robustness dimension is not evaluated herein. trustworthiness index Ranking of synthetic data sampled from different epoch during the finetuning of BioGPT model (Note that the robustness dimension was not considered) . The ranking corresponds to different trust tradeoffs $\omega$: for $w_{U}$ that is accuracy driven, we see that the last epoch is outperforming the other ones; When in addition we consider the fairness of the prediction ($\omega_{UF}$) the last epoch (epoch 9) ranks third and the epoch 7 presents better utility/ Fairness trade-offs.
Training RoBERTa-like Embedding
To train TabRoBERTa on our transaction dataset, we constructed samples as sliding windows of 10 transactions, using a stride of 5. We excluded the label column, "isFraud?", during training to prevent biasing the learned representation for the downstream fraud detection task. We masked $15\%$ of a sample’s fields, replacing them with the [MASK] token, and predicted the original field token using cross-entropy loss. We used DP-SGD for transformer models [68] to train various RoBERTa-like models with differing degrees of privacy, ranging from highly private ($\varepsilon=1$) to non-private ($\varepsilon=1000$). Additionally, we trained a RoBERTa model without private training (see the columns labeled "Private" and "Non-Private" in Table VII).